# Consistency of JWST Black Hole Observations with NANOGrav Gravitational Wave Measurements

### v1

@@ Abstract
JWST observations have opened a new chapter in studies of supermassive black holes (SMBHs), stimulating discussion of two puzzles: the abundance of SMBHs in the early Universe and the fraction of dual AGNs. In this paper we argue that the answers to these puzzles may be linked to an interpretation of the data on the nHz gravitational wave (GWs) discovered by NANOGrav and other Pulsar Timing Arrays (PTAs) in terms of SMBH binaries losing energy by interactions with their environments as well as by GW emission. According to this interpretation, the SMBHs in low-$z$ AGNs are the tip of the iceberg of the local SMBH population, which are mainly in inactive galaxies. This interpretation would favour the observability of GW signals from BH binaries in LISA and deciHz GW detectors.
@@
Introduction – There has been some surprise at the number of supermassive black holes (SMBHs) discovered with the JWST at high redshifts $z$: see Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. ; Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023), the red data points shown in Fig. 1 and the scatter plot in the Appendix. Prior expectations for the number of SMBHs in the early Universe were largely based on empirical relationships between the SMBH mass-halo mass ratio and the SMBH mass-stellar mass ratio, such as that found by Reines and Volonteri (RV) in the local (low-$z$) population of active galactic nuclei (AGNs) Reines and Volonteri (2015): see the green points in Fig. 1. In a statistical analysis, Pacucci et al. Pacucci et al. (2023) concluded that there was a $>3-\sigma$ discrepancy between the SMBH mass/stellar mass ratios, $M_{\rm BH}/M_{*}$, of a subset of the high-$z$ JWST data and this fit to local AGN data. We note, however, that RV also analysed a sample of local Inactive Galaxies (IGs), finding higher values of $M_{\rm BH}/M_{*}$ Reines and Volonteri (2015): see the blue points in Fig. 1. Moreover, larger values of $M_{\rm BH}/M_{*}$ had also been found in a compilation of high-$z$ SMBHs observed previously to JWST in Izumi et al. (2021): see the yellow data in Fig. 1.
@@
Another surprise in JWST data was the discovery of several close pairs in a sample of AGNs Perna et al. (2023), whereas cosmological simulations generally predicted a fraction of such dual AGNs that was at least an order of magnitude smaller.
@@
In parallel with the emergence of JWST data, Pulsar Timing Arrays (PTAs) have reported the existence of a stochastic gravitational wave background (SGWB) in the nHz range Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), whose most plausible astrophysical interpretation is gravitational wave (GW) emission by SMBH binaries. Analysing their data under this assumption and guided by astrophysical priors, the NANOGrav Collaboration found Agazie et al. (2023b) that their data were best fit by a $M_{\rm BH}/M_{*}$ ratio that was significantly higher than that found in the RV analysis of local AGNs Reines and Volonteri (2015). In parallel, we found independently in Ellis et al. (2024a), using the Extended Press-Schechter formalism Press and Schechter (1974); Bond et al. (1991) to estimate the halo merger rate and hence the evolution of the halo mass function, and assuming a fixed probability, $p_{\rm BH}$, for a halo merger to yield a BH merger, that a good fit to the NANOGrav data was possible with the $M_{\rm BH}/M_{*}$ ratio found by RV in their analysis of local IGs Reines and Volonteri (2015). It was also found in  Ellis et al. (2024a) that using the RV fit to local AGNs degraded significantly the fit to the NANOGrav data: $\Delta\chi^{2}=43$ compared to the fit based on local IGs: see Fig. 9 of Ellis et al. (2024a) and the discussion in the accompanying text, as well as the analysis below.
@@
It has been suggested recently that the $M_{\rm BH}/M_{\rm*}$ relationship may evolve with $z$ Pacucci and Loeb (2024), or that selection biases and measurement uncertainties Li et al. (2024) may account for much of the tension between the JWST high-$z$ SMBH data and the data on local AGNs considered in Reines and Volonteri (2015). However, we take at face value the prima facie evidence for consistency between the information on SMBHs provided by the high-$z$ SMBH data and NANOGrav measurements. The puzzle of the masses of JWST black holes is apparently linked to the somewhat unexpected strength of the GW signal reported by NANOGrav and other PTAs Agazie et al. (2023c), and can be reframed as a puzzle why both data sets correspond better to the data on local IGs than to local AGNs.
@@
In this paper we quantify in more detail the consistency between the NANOGrav nHz GW data and the JWST and previous data on high-$z$ SMBHs, discussing both the SMBH mass/stellar mass ratio and the fraction of dual AGNs in the JWST data. We confirm the tension between the SMBH mass-stellar mass ratio found in the population of low-$z$ AGNs and a subset of the JWST data found in Pacucci et al. (2023). However, we also show that the JWST data are much more compatible with the low-$z$ IG data, and that the JWST data are also compatible at an even higher level with the compilation of earlier SMBH data in Izumi et al. (2021). Moreover, we show that the high-$z$ SMBH data are also consistent with our analysis of the NANOGrav data on the nHz GW background Ellis et al. (2024a). Our interpretation of the NANOGrav and JWST data would favour the observability of GW signals from BH binaries in LISA and deciHz GW detectors, as discussed previously in Ellis et al. (2024a). It can also accommodate the unexpectedly high fraction of dual AGNs observed with JWST Perna et al. (2023). Assuming that the hardening timescale due to halo-halo interactions is similar to that found empirically for environmental effects in our fit to NANOGrav data Ellis et al. (2024a), we again find consistency at the 68% confidence level, suggesting linkage also to the dual AGN puzzle.
@@
Figure 1: Measurements and fits of the black hole mass–stellar mass relation. The data points with errors are from JWST measurements (red, see the Appendix) and other high-$z$ measurements (yellow) Izumi et al. (2021). The local AGN measurements (green) and the local IG measurements (blue), both from Reines and Volonteri (2015), are shown without errors, but the errors are taken into account in the analysis. The dashed lines show the best fits to each of the data sets separately and the black solid line shows the fit combining the high-$z$ data and the local IGs data. The gray band shows the scatter in the latter fit.
@@
As commonly done in the literature, we assume a power-law form for the relation between the SMBH mass and the stellar mass of the host galaxy, with a log-normal intrinsic scatter. The probability distribution of the SMBH mass $m_{\rm BH}$ for a given stellar mass $M_{*}$ is then parametrized by the magnitude $a$ and logarithmic slope $b$ of the mean and the width $\sigma$ of the distribution:
$\frac{{\rm d}P(M_{*},\boldsymbol{\theta})}{{\rm d}\log_{10}\!m_{\rm BH}}=\mathcal{N}\bigg{(}\!\log_{10}\!\frac{m_{\rm BH}}{M_{\odot}}\bigg{|}a+b\log_{10}\!\frac{M_{*}}{10^{11}M_{\odot}},\sigma\bigg{)}\,,$ (1)
@@
$\mathcal{L}(\boldsymbol{\theta})$ $\propto\prod_{j}\int\!{\rm d}\log_{10}\!m\,{\rm d}\log_{10}\!M\,\frac{{\rm d}P(M,\boldsymbol{\theta})}{{\rm d}\log_{10}m}$ (2)
@@
where the product is over the data points, $m_{j}$ and $M_{j}$ denote the mean measured BH and stellar masses, and we assume that posteriors of the BH and stellar mass measurements are log-normal and uncorrelated. We use flat priors for $a$, $b$ and $\sigma$ with $b>0$.
As had been seen in Fig. 1, the high-$z$ AGN measurements are largely consistent with each other and with the fit to the local IGs, though the latter prefers slightly stronger tilt. The posteriors in $a$ and $b$ found in a global fit to these data, marginalized over $\sigma$, are shown in the upper panel of Fig. 2, and the posteriors of the fits to individual data sets are shown in the Appendix. The lower panel of Fig. 2 displays uncertainty ellipses for the individual data sets as well as that for the global fit (shown in gray) that omits the local AGN data (shown in green). The global best fit is given by
@@
Figure 2: Upper Panel: Corner plot showing the posteriors for the global fit to JWST Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. ; Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023) and previous high-$z$ SMBH data Izumi et al. (2021) as well as low-$z$ data on SMBHs in inactive galaxies Reines and Volonteri (2015). Lower Panel: The 68%, 95% and 99.8% CL contours (gray) of the posteriors for $a$ and $b$ defined in (1), obtained in a global fit including all observations except local AGNs by marginalizing over $\sigma$. For comparison, the 68% and 99.8% CL of individual data sets are shown: JWST high-$z$ measurements (red) Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. ; Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023), other high-$z$ measurements (yellow) Izumi et al. (2021), local inactive galaxies (blue) and local AGNs (green) Reines and Volonteri (2015). The fits are performed using flat priors and including the uncertainties in the observations. Corner plots showing the posterior distributions for these fits are displayed in the Appendix.
@@
In Ellis et al. (2024a) we found that the local IG fit also gives a much better fit to the NANOGrav GW observations than does the local AGN fit, and we now develop this point further. As in Ellis et al. (2023, 2024a, 2024b), we relate the SMBH merger rate to the halo merger rate $R_{h}$, estimated from the Press-Schechter formalism (Press and Schechter, 1974; Bond et al., 1991; Lacey and Cole, 1993), as
$\frac{{\rm d}R_{\rm BH}(p_{\rm BH},\boldsymbol{\theta})}{{\rm d}m_{1}{\rm d}m_{2}}\approx$ $\,\,p_{\rm BH}\int{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}\frac{{\rm d}R_{h}}{{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}}$ (5)
×$\hskip 10.0pt\times\prod_{j=1,2}\frac{{\rm d}P(M_{*}(M_{{\rm v},j},z),\boldsymbol{\theta})}{{\rm d}m_{j}}\,,$
where $p_{\rm BH}$ characterizes the BH merger efficiency and $M_{*}(M_{{\rm v},j},z)$ is the halo mass-stellar mass relation that we take from Girelli et al. (2020). We perform the analysis of the NANOGrav GW signal in the same way as in Ellis et al. (2024a) including the environmental effects parametrized by a spectral index $\alpha$ and a reference frequency $f_{\rm ref}$ (see also Eq. (11)). Fig. 3 displays in a corner plot the posterior distributions for the parameters of this fit: $\alpha,f_{\rm ref}$ and $p_{\rm BH}$. The posteriors for the parameters $\alpha$ and $f_{\rm ref}$ are compatible with models of environmental effects, and the preferred range of the posterior for $p_{\rm BH}\in[0.55,1]$.
@@
In the upper panel of Fig. 4 we compare our global fit (3) with the NANOGrav SGWB data Agazie et al. (2023a) and the best fit SMBH binary GW signals obtained with the BH mass-stellar mass relations derived in Pacucci et al. (2023) and Li et al. (2024). The data and the best fit simulations are represented by “violins” whose widths represent the posteriors as functions of the GW density: as in Ellis et al. (2024a), we calculate the overlaps between the violins to assess the qualities of the fits. For both Pacucci et al. (2023); Li et al. (2024) the generated SGWB is insufficient and the best fit saturates at $p_{\rm BH}=1$, which means that every galactic merger leads to a GW signal. We see that the fits are particularly poor at frequencies below 10 nHz. On the other hand, in our analysis with the global fit mass relation (3), the best fit is achieved at $p_{\rm BH}=0.8$, $f_{\rm ref}=15\,$nHz and $\alpha=2$. By comparing the best-fit likelihoods, we find that the PTA GW background favours the global fit mass relation at higher than $5\sigma$ CL compared to the scaling relations proposed in Pacucci et al. (2023); Li et al. (2024).
@@
We show in the lower panel of Fig. 4 50 random realizations of the strongest binary contributing to the first NANOGrav bin, $f_{\rm GW}\in(2-4)$ nHz, for each of the three scenarios. In the case of Li et al. (2024), which advocated for a strong bias in the scaling relation, we can see that the SMBHs are too light, whereas in the case of Pacucci and Loeb (2024), which advocates for a steep $z$ dependence, the heavy SMBH are merging mainly at $z>1$, which also decreases the amplitude significantly.
@@
Figure 4: Upper panel: Best fit nHz GW signals found using the scaling relations proposed in Pacucci and Loeb (2024) and Li et al. (2024) compared to the signal measured by NANOGrav (gray) and our analysis Ellis et al. (2024a) with the global fit mass relation (3). Lower panel: Chirp masses and redshifts of 50 realizations of the strongest binary contributing to the first bin of the upper panel for the scenarios proposed in Pacucci and Loeb (2024); Li et al. (2024), compared with our analysis Ellis et al. (2024a).
@@
Furthermore, in the lower panel of Fig. 4 we also see that black holes with $m_{\rm BH}<10^{9}\,M_{\odot}$ are very unlikely to be among the strongest sources contributing to the SGWB. Given that the mass and $z$ range probed by NANOGrav coincides with the dynamically measured SMBH masses, the high merger efficiency necessary to explain the observed GW background implies that the local IG population is not significantly biased. Since the local AGNs lie in a lower mass range, it is an open question whether that population is biased. However, given the consistency between the high-$z$ AGN measurements and the local IG measurements and the fact that at high $z$ the AGN fraction is close to unity but drops quickly at $z\lesssim 1$ Aird et al. (2018); Georgakakis et al. (2017), it seems that the local AGN population reflects the intrinsic low-mass tail of the full BH mass-stellar mass relationship.
@@
Comparing the local AGN population to the global fit in Fig. 1 we see that the majority of the local AGNs are more than $2\sigma$ below the mean of the global fit. Hence, according to this analysis, the AGN fraction in the local SMBH population can only be $\lesssim 5\%$, in agreement with X-ray observations Aird et al. (2018); Georgakakis et al. (2017). To conclude, in contrast to Li et al. (2024), our analysis indicates that the local AGN population represents the tip of the iceberg of the local SMBH population. As discussed in Ellis et al. (2024a, c), our interpretation of the NANOGrav and JWST data would favor the observability of GW signals from BH binaries in space-borne laser interferometers (e.g., LISA Amaro-Seoane et al. (2017), TianQin Wang et al. (2019) and Taiji Ruan et al. (2020)) and deciHz GW detectors (e.g., AION Badurina et al. (2020), AEDGE El-Neaj et al. (2020) and DECIGO Kawamura et al. (2021)).
@@
At the kpc separations dynamical friction drives the binary evolution Kelley et al. (2017). Its timescale is Binney and Tremaine (2008)
@@
where $\ln\Lambda$ is the Coulomb logarithm, $\sigma$ is the velocity dispersion of stars and $M$ is the binary total mass. However, dynamical friction becomes inefficient at shorter distances when the binary binding energy becomes larger than the kinetic energy of the stars whose orbits pass close to the binary. The binary evolution is then driven by loss-cone scattering, whose timescale can be estimated as Quinlan (1996)
$t_{\rm sh}\simeq\frac{\sigma}{H\rho d}\approx 300\,{\rm Myr}\frac{\sigma}{200\,{\rm km}/s}\!\left[\frac{\rho}{10M_{\odot}/{\rm pc}^{3}}\right]^{-1}\!\left[\frac{d}{\rm pc}\right]^{-1}\!\!,$ (10)
where $H\approx 15-20$ is the dimensionless hardening rate and $\rho$ and $\sigma$ denote the mass density and the velocity dispersion of stars in the neighborhood of the binary.
At the intermediate scales between the GW-driven and the loss-cone scattering-driven phases, we follow Ellis et al. (2024a) by parametrizing the environmental effects with a power-law

### v2

@@ Abstract
JWST observations have opened a new chapter in supermassive black hole (SMBH) studies, stimulating discussion of two puzzles: the abundance of SMBHs in the early Universe and the fraction of dual AGNs. In this paper, we argue that the answers to these puzzles may be linked to an interpretation of the data on the nHz gravitational wave (GW) discovered by NANOGrav and other pulsar timing arrays (PTAs) in terms of SMBH binaries losing energy by interactions with their environments as well as by GW emission. According to this interpretation, the SMBHs in low-$z$ AGNs are the tip of the iceberg of a local SMBH population populating mainly inactive galaxies. This interpretation is consistent with the “little red dots” observed with JWST being AGNs, and would favour the observability of GW signals from BH binaries in LISA and deciHz GW detectors.
@@
Introduction – There has been some surprise at the number of supermassive black holes (SMBHs) discovered with the JWST at high redshifts $z$: see Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. (2024); Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023), the red data points shown in Fig. 1 and the scatter plot in the Appendix. Prior expectations for the number of SMBHs in the early Universe were largely based on empirical relationships between the SMBH mass-halo mass ratio and the SMBH mass-stellar mass ratio, such as that found by Reines and Volonteri (RV) in the local (low-$z$) population of active galactic nuclei (AGNs) Reines and Volonteri (2015): see the green points in Fig. 1. In a statistical analysis, Pacucci et al. Pacucci et al. (2023) concluded that there was a $>3\sigma$ discrepancy between the SMBH mass/stellar mass ratios, $M_{\rm BH}/M_{*}$, of a subset of the high-$z$ JWST data and this fit to local AGN data. We note, however, that RV also analysed a sample of local inactive galaxies (IGs), finding higher values of $M_{\rm BH}/M_{*}$ Reines and Volonteri (2015): see the blue points in Fig. 1. Moreover, larger values of $M_{\rm BH}/M_{*}$ had also been found in a compilation of high-$z$ SMBHs observed previously by JWST in Izumi et al. (2021): see the yellow data in Fig. 1.
@@
Another surprise in JWST data was the discovery of several close pairs in a sample of AGNs Perna et al. (2023), whereas cosmological simulations generally predicted a fraction of such dual AGNs that was at least an order of magnitude smaller. This observation was used in Padmanabhan and Loeb (2024) to study SMBH binary evolution. JWST observations have also revealed a population of “little red dots” (LRDs) in the redshift range $z\sim 2-11$ that are interpreted as previously-hidden AGNs tracing BH growth in the early Universe Labbé et al. (2023a); Akins et al. (2023); Barro et al. (2024); Labbé et al. (2023b); Matthee et al. (2024); Kocevski et al. (2024).
@@
In parallel with the emergence of JWST data, pulsar timing arrays (PTAs) have reported the existence of a stochastic gravitational wave background (SGWB) in the nHz range Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), whose most plausible astrophysical interpretation is gravitational wave (GW) emission by SMBH binaries. Analyzing their data under this assumption and guided by astrophysical priors, the NANOGrav Collaboration found Agazie et al. (2023b) that their data were best fit by a $M_{\rm BH}/M_{*}$ ratio that was significantly higher than that found in the RV analysis of local AGNs Reines and Volonteri (2015). In parallel, we found independently in Ellis et al. (2024a), using the Extended Press-Schechter formalism Press and Schechter (1974); Bond et al. (1991) to estimate the halo merger rate and hence the evolution of the halo mass function, and assuming a fixed probability, $p_{\rm BH}$, for a halo merger to yield a BH merger, that a good fit to the NANOGrav data was possible with the $M_{\rm BH}/M_{*}$ ratio found by RV in their analysis of local IGs Reines and Volonteri (2015). It was also found in  Ellis et al. (2024a) that using the RV fit to local AGNs degraded significantly the fit to the NANOGrav data: $\Delta\chi^{2}=43$ compared to the fit based on local IGs: see Fig. 9 of Ellis et al. (2024a) and the discussion in the accompanying text, as well as the analysis below.
@@
It has been suggested recently that the $M_{\rm BH}/M_{\rm*}$ relationship may evolve with $z$ Pacucci and Loeb (2024), or that selection biases and measurement uncertainties Li et al. (2024) may account for much of the tension between the JWST high-$z$ SMBH data and the data on local AGNs considered in Reines and Volonteri (2015). However, we take at face value the prima facie evidence for consistency between the information on SMBHs provided by the high-$z$ SMBH data and NANOGrav measurements. The puzzle of the masses of JWST black holes may be linked to the somewhat unexpected strength of the GW signal reported by NANOGrav and other PTAs Agazie et al. (2023c), and can be reframed as a puzzle why both data sets correspond better to the data on local IGs than to local AGNs.
@@
In this paper we quantify in more detail the consistency between the NANOGrav nHz GW data and the JWST and previous data on high-$z$ SMBHs, discussing both the SMBH mass/stellar mass ratio and the fraction of dual AGNs in the JWST data. We confirm the tension between the SMBH mass-stellar mass ratio found in the population of low-$z$ AGNs and a subset of the JWST data found in Pacucci et al. (2023). However, we also show that the JWST data are much more compatible with the low-$z$ IG data, and that the JWST data are also compatible at an even higher level with the compilation of earlier SMBH data in Izumi et al. (2021). Moreover, we show that the high-$z$ SMBH data are also consistent with our analysis of the NANOGrav data on the nHz GW background Ellis et al. (2024a). Our interpretation of the NANOGrav and JWST data would favour the observability of GW signals from BH binaries in LISA and deciHz GW detectors, as discussed previously in Ellis et al. (2024a). It can also accommodate the unexpectedly high fraction of dual AGNs observed with JWST Perna et al. (2023). Assuming that the hardening timescale due to halo-halo interactions is similar to that found empirically for environmental effects in our fit to NANOGrav data Ellis et al. (2024a), we find consistency at the 95% confidence level, suggesting linkage also to the dual AGN puzzle– which also favours the observability of GWs from massive BH mergers. In addition, our interpretation of the JWST and NANOGrav data is consistent with the population of “little red dots” reported recently Matthee et al. (2024); Kocevski et al. (2024), assuming that they are AGNs triggered during mergers as discussed in more detail below.
@@
Figure 1: Measurements and fits of the black hole mass–stellar mass relation. The data points with errors are from JWST measurements (red, see the Appendix) and other high-$z$ measurements (yellow) Izumi et al. (2021). The local AGN measurements (green) and the local IG measurements (blue), both from Reines and Volonteri (2015), are shown without errors, but the errors are taken into account in the analysis. The dashed lines show the best fits to each of the data sets separately and the black solid line shows the best fit combining the high-$z$ data and the local IGs data. The grey band shows the scatter in the latter fit.
@@
As commonly done in the literature, we assume a power-law form for the relation between the SMBH mass and the stellar mass of the host galaxy, with a log-normal intrinsic scatter. The probability distribution of the SMBH mass $m$ for a given stellar mass $M_{*}$ is then parametrized by the magnitude $a$ and logarithmic slope $b$ of the mean and the width $\sigma$ of the distribution:
$\frac{{\rm d}P(m|M_{*},\boldsymbol{\theta})}{{\rm d}\log_{10}\!m}=\mathcal{N}\bigg{(}\!\log_{10}\!\frac{m}{M_{\odot}}\bigg{|}a+b\log_{10}\!\frac{M_{*}}{10^{11}M_{\odot}},\sigma\bigg{)}\,,$ (1)
@@
$\mathcal{L}(\boldsymbol{\theta})$ $\propto\prod_{j}\int\!{\rm d}\log_{10}\!m\,{\rm d}\log_{10}\!M\,\frac{{\rm d}P(m|M,\boldsymbol{\theta})}{{\rm d}\log_{10}m}$ (2)
@@
where the product is over the data points, $m_{j}$ and $M_{j}$ denote the mean measured BH and stellar masses, and we assume that the posteriors of the BH and stellar mass measurements are log-normal and uncorrelated. We use flat priors for $a$, $b$ and $\sigma$ with $b>0$.
As seen in Fig. 1, the high-$z$ AGN measurements are largely consistent with each other and with the fit to the local IGs, though the latter prefers a slightly stronger tilt. The posteriors in $a$ and $b$ found in a global fit to these data, marginalized over $\sigma$, are shown in the upper panel of Fig. 2, and the posteriors of the fits to individual data sets are shown in the Appendix. The lower panel of Fig. 2 displays uncertainty ellipses for the individual data sets as well as that for the global fit (shown in grey) that omits the local AGN data (shown in green). The global best fit is given by
@@
Figure 2: Upper Panel: Corner plot showing the posteriors for the global fit to JWST Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. (2024); Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023) and previous high-$z$ SMBH data Izumi et al. (2021) as well as low-$z$ data on SMBHs in IGs Reines and Volonteri (2015). Lower Panel: The 68%, 95% and 99.8% CL contours (gray) of the posteriors for $a$ and $b$ defined in (1), obtained in a global fit including all observations except local AGNs by marginalizing over $\sigma$. For comparison, the 68% and 99.8% CL of individual data sets are shown: JWST high-$z$ measurements (red) Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. (2024); Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023), other high-$z$ measurements (yellow) Izumi et al. (2021), local IGs (blue) and local AGNs (green) Reines and Volonteri (2015). The fits are performed using flat priors and including the uncertainties in the observations. Corner plots showing the posterior distributions for these fits are displayed in the Appendix.
@@
In Ellis et al. (2024a) we found that the local IG fit also gives a much better fit to the NANOGrav GW observations than the local AGN fit, and we now develop this point further. As in Ellis et al. (2023, 2024a, 2024b), we relate the SMBH merger rate to the halo merger rate $R_{h}$, estimated from the Press-Schechter formalism (Press and Schechter, 1974; Bond et al., 1991; Lacey and Cole, 1993), as
$\frac{{\rm d}R_{\rm BH}}{{\rm d}m_{1}{\rm d}m_{2}}\approx$ $\,\,p_{\rm BH}\int{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}\frac{{\rm d}R_{h}}{{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}}$ (5)
×$\hskip 10.0pt\times\prod_{j=1,2}\frac{{\rm d}P(m_{j}|M_{*}(M_{{\rm v},j},z))}{{\rm d}m_{j}}\,,$
where $p_{\rm BH}$ characterizes the BH merger efficiency and $M_{*}(M_{{\rm v},j},z)$ is the halo mass-stellar mass relation that we take from Girelli et al. (2020). We perform the analysis of the NANOGrav GW signal in the same way as in Ellis et al. (2024a), including environmental effects parametrized by a spectral index $\alpha$ and a reference frequency $f_{\rm ref}$ (see Eq. (11) and the surrounding discussion). Fig. 3 displays in a corner plot the posterior distributions for the parameters of this fit: $\alpha,f_{\rm ref}$ and $p_{\rm BH}$. The posterior for $\alpha$ is compatible with models of environmental effects and the fit prefers large values of $f_{\rm ref}\gtrsim 10$ nHz. The preferred range of $p_{\rm BH}$ is $[0.16,0.38]$. The best fit is achieved at $p_{\rm BH}=0.37$, $f_{\rm ref}=27\,$nHz and $\alpha=3.3$. In particular, the value $\alpha=8/3$ expected for gas infall agrees well with the data. By fixing $\alpha=8/3$, we find that the best fit is given by $p_{\rm BH}=0.37$, $f_{\rm ref}=30\,$nHz.
@@
In the upper panel of Fig. 4 we compare our global fit (3) with the NANOGrav SGWB data Agazie et al. (2023a) and the best fit SMBH binary GW signals obtained with the BH mass-stellar mass relations derived in Pacucci et al. (2023) and Li et al. (2024). The data and the best-fit simulations are represented by “violins” whose widths represent the posteriors as functions of the GW density: following Ellis et al. (2024a), we calculate the overlaps between the violins to assess the qualities of the fits. For both Pacucci et al. (2023) and Li et al. (2024) the best fits saturate at $p_{\rm BH}=1$, i.e., every galactic merger leads to a GW signal, which is an indication that the generated SGWB is insufficient. Comparing the best-fit likelihoods, we find that the PTA GW background favours the global fit mass relation at higher than the $3\sigma$ CL compared to the scaling relations proposed in Pacucci et al. (2023); Li et al. (2024).
@@
We show in the lower panel of Fig. 4 100 random realizations of the strongest binary contributing to the first NANOGrav bin, $f_{\rm GW}\in(2-4)$ nHz, for each of the three scenarios. In the case of Li et al. (2024), which advocated for a strong bias in the scaling relation, we can see that the SMBHs are too light, whereas in the case of Pacucci and Loeb (2024), which advocates for a steep $z$ dependence, the heavy SMBH are merging mainly at $z>1$, which also decreases the amplitude significantly.
@@
Figure 4: Upper panel: Best fit nHz GW signals found using the BH mass-stellar mass relations proposed in Pacucci and Loeb (2024), Li et al. (2024) and our global fit (3) compared to the signal measured by NANOGrav (gray). We imposed $\alpha=8/3$ as expected from gas infall (see discussion below Eq. (11)). Lower panel: Chirp masses and redshifts of 100 realizations of the strongest binary contributing to the first bin of the upper panel for the scenarios proposed in Pacucci and Loeb (2024); Li et al. (2024), compared with our analysis Ellis et al. (2024a). Contours of $\log_{10}\Omega_{\rm GW}$ are shown for comparison.
@@
Furthermore, in the lower panel of Fig. 4 we also see that black holes with $m_{\rm BH}<10^{9}\,M_{\odot}$ are very unlikely to be among the strongest sources contributing to the SGWB. Given that the mass and $z$ range probed by NANOGrav coincides with the dynamically measured SMBH masses, the high merger efficiency necessary to explain the observed GW background implies that the local IG population is not significantly biased. The local AGNs lie in a lower mass range, and it is an open question whether that population is biased. However, given the consistency between the high-$z$ AGN measurements and the local IG measurements and the fact that at high $z$ the AGN fraction is close to unity but drops quickly at $z\lesssim 1$ Aird et al. (2018); Georgakakis et al. (2017), it seems that the local AGN population reflects the intrinsic low-mass tail of the full BH mass-stellar mass relationship.
@@
Comparing the local AGN population to the global fit in Fig. 1 we see that most of the local AGNs lie more than $2\sigma$ below the mean of the global fit. Hence, according to this analysis, the AGN fraction in the local SMBH population can only be $\lesssim 5\%$, in agreement with X-ray observations Aird et al. (2018); Georgakakis et al. (2017). To conclude, in contrast to Li et al. (2024), our analysis indicates that the local AGN population represents the tip of the iceberg of the local SMBH population. As discussed in Ellis et al. (2024a, c), our interpretation of the NANOGrav and JWST data would favour the observability of GW signals from BH binaries in space-borne laser interferometers (e.g., LISA Amaro-Seoane et al. (2017), TianQin Wang et al. (2019) and Taiji Ruan et al. (2020)) and deciHz GW detectors (e.g., AION Badurina et al. (2020), AEDGE El-Neaj et al. (2020) and DECIGO Kawamura et al. (2021)).
@@
At kpc separations, dynamical friction drives the binary evolution Kelley et al. (2017). Its timescale is Binney and Tremaine (2008)
@@
where $\ln\Lambda$ is the Coulomb logarithm, $\sigma$ is the velocity dispersion of stars.
Dynamical friction becomes inefficient at shorter distances when the binary binding energy becomes larger than the kinetic energy of the stars whose orbits pass close to the binary. The binary evolution is then driven by loss-cone scattering, whose timescale can be estimated as Quinlan (1996)
$t_{\rm sh}\simeq\frac{\sigma}{H\rho d}\approx 300\,{\rm Myr}\frac{\sigma}{200\,{\rm km}/{\rm s}}\!\left[\frac{\rho}{10M_{\odot}/{\rm pc}^{3}}\right]^{-1}\!\left[\frac{d}{\rm pc}\right]^{-1}\!\!,$ (10)
where $H\approx 15-20$ is the dimensionless hardening rate and $\rho$ and $\sigma$ denote the mass density and the velocity dispersion of stars in the neighbourhood of the binary.
At intermediate scales between the GW-driven and loss-cone scattering-driven phases, as mentioned earlier we follow Ellis et al. (2024a) by parametrizing the environmental effects with a power-law